Factorise the following using appropriate identities:
:l
9r3 + 6zy + i' cu> 4y2 - 4y + 1 (iii> r2 - L.
100

Grade:12th pass

1 Answers

Pawan Prajapati
3 years ago
(i) 9%2 + &:r:,i + • C&.>2 + 2 x & x y + (y)2 = (3% + y f [Using identity:x2 + 2xy + y2 = (% + y)2J 38 = (3x + y)(3x + y) (ii) - 4y + 1= (2y)2 - 2 )( 2y )( 1+ (1)2 = (2y - 1)2 [Using identity:t:2 - 2%y + y2 = y)2] = (2y - 1)(2y - 1). (iii) r2 _ L = (x)2 _ (Ll= (x-L) (x +L). [Using identity: a2 - b2 = (a - bXa + b)]